OpenStudy (girlgoyleh.):

Eric has a gift card to his favorite restaurant. Every Tuesday he has the same meal at the restaurant. What is the equation that represents Eric's gift card balance? Record any fractional answers in decimal form. (y -100) =___(x - ___ ) Eric's Gift Card Weeks 2 4 6 8 Balance$125 $100 $75 $50

OpenStudy (triciaal):

you can calculate the change or plot the graph and see the change is $-75/2 weeks using the point slope formula the point in the equation has a y value of 100

OpenStudy (triciaal):

m = slope = change in y/ corresponding change in x m = (Y2 - y1)/(x2 - x1) using 2 points (x1,y1) and (x2, y2) on that line

OpenStudy (triciaal):

* not -75 -25

OpenStudy (texaschic101):

find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) using 2 of your points....any 2 (2,125)...x1 = 2 and y1 = 125 (4,100)...x2 = 4 and y2 = 100 now sub into the formula to find the slope

OpenStudy (texaschic101):

let me get you started.. slope(m) = (100 - 125) / (4 - 2) can you find the slope now ?

OpenStudy (girlgoyleh.):

that / is the division symbol?

OpenStudy (texaschic101):

yes..but leave it as a fraction

OpenStudy (girlgoyleh.):

-25/2

OpenStudy (texaschic101):

correct...so that is your slope. Now we will use point-slope formula y - y1 = m(x - x1)

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = -25/2 It looks like your answer already has 100 as y....so we will use points (4,100)..x1 = 4 and y1 = 100. Now we sub y - 100 = -25/2(x - 4) <== your answer any questions ?

OpenStudy (girlgoyleh.):

thank you so much! that was surprisingly easy...

OpenStudy (texaschic101):

it is not as hard as it looks...lol glad I could help :)

OpenStudy (girlgoyleh.):

i got it wrong, it was 12.5...

OpenStudy (triciaal):

@GirlgoyleH do you understand why you got it wrong?

OpenStudy (triciaal):

you did not follow the directions. Record any fractional answers in decimal form. -25/2 = negative 12.5

OpenStudy (chunkymonkay):

@triciaal so the answer would be (y-100)=12.5(x-4)?

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

dont forget the - on 12.5